OpenStudy (anonymous):

How to differentiate this: y = (x^3 - 3x^2 + 1)^9

OpenStudy (anonymous):

product and chain rules

OpenStudy (freckles):

\[\frac{d}{dz}(z)^9|_{z=x^3-3x^2+1} \cdot (x^3-3x^2+1)'\]

OpenStudy (anonymous):

y = 9(x^3 - 3x^2 + 1)^8 * (x^3 - 3x^2 + 1)

OpenStudy (anonymous):

?

OpenStudy (freckles):

Well you got the outside derivative right but you still didn't take derivative of the inside

OpenStudy (freckles):

\[y'=\frac{d}{dz}(z)^9|_{z=x^3-3x^2+1} \cdot (x^3-3x^2+1)' \\ y'=9(x^3-3x^2+1)^8 \cdot (x^3-3x^2+1)'\]

OpenStudy (anonymous):

Okay. 9(3x^2 - 6x)(x^3 - 3x^2 +1)^8

OpenStudy (freckles):

looks tons better

OpenStudy (anonymous):

Then do we do this:

OpenStudy (anonymous):

a = (3x^2 - 6x) b = (x^3 - 3x^2 +1)^8 a^1 x b - a x b^1

OpenStudy (freckles):

when i said tons better means it was good it was right

OpenStudy (freckles):

what are you doing now?

OpenStudy (anonymous):

oh okay. nevermind.

OpenStudy (anonymous):

thanks. :)

OpenStudy (freckles):

unless you want to find the second derivative ?

OpenStudy (freckles):

because when I look at it I think that is what you want to do now since you were applying product rule and we have a product and I guess by f^1 that is f'